# Asymptotic behaviour of the Hitchin metric on the moduli space of Higgs bundles

### v2

@@ 3.1.1 Spectral curves
We recall the spectral curves by following [1]. Let $X$ be a compact Riemann surface. In this section, we prefer to use the notation $K_{X}$ to denote the sheaf of holomorphic $1$-forms because we consider $K_{X}^{j/2}$. We set
@@
Let $Z_{s}$ be the subscheme of $\mathbb{P}$ obtained as the 00-set of $\xi_{s}$. It is contained in $T^{\ast}X\subset\mathbb{P}$. Let ${\boldsymbol{A}}^{\prime}_{X,n}$ denote the subset of $s\in{\boldsymbol{A}}^{\prime}_{X,n}$ such that $Z_{s}$ are smooth. Let $\pi_{s}:Z_{s}\to X$ denote the projection. For each $Q\in Z_{s}$, let $r_{Q}$ denote the ramification index of $\pi_{s}$ at $Q$. We obtain the effective divisor $\widetilde{D}_{s}=\sum(r_{Q}-1)Q$. We have $K_{Z_{s}}=\pi^{\ast}(K_{X})(\widetilde{D}_{s})$. We recall the following.
@@
Let $g_{X}$ denote the genus of $X$.
  • For any $s\in{\boldsymbol{A}}^{\prime}_{X,n}$, the genus of $Z_{s}$ is $n^{2}(g_{X}-1)+1$.
  • $\deg(\widetilde{D}_{s})=2n(n-1)(g_{X}-1)$.
@@ 3.1.2 Singular flat metrics
Let $L_{i}$ $(i=1,2)$ be holomorphic line bundles on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. Let $U_{j}$ $(j=1,\ldots,m)$ be simply connected open subsets of $X$ such that $U_{j}\cap U_{k}=\emptyset$ $(j\neq k)$ and that $|\widetilde{D}_{s}|\subset\bigcup U_{j}$. Let
@@ 3.1.3 A special case
The image of $\kappa_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}:H_{1}(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}})\to U(1)$ is contained in {±$\{\pm 1\}$.
Proof   Recall that $Z_{s}$ is the 00-set of $\xi_{s}\in H^{0}(\mathbb{P},\pi_{\mathbb{P}}^{\ast}K_{X}^{n}\otimes\mathcal{O}_{\mathbb{P}}(n))$, and that the line bundle $\mathcal{O}_{\mathbb{P}}(n)$ is trivialized by the section $\pi_{\mathbb{P}}^{\ast}(y)$ on a neighbourhood of $Z_{s}$. (See §3.1.1 for the section $y$.) Hence, there exists a natural isomorphism between $\pi_{s}^{\ast}K_{X}^{n}$ and the normal bundle of $Z_{s}\subset T^{\ast}X$. By the symplectic structure of $T^{\ast}X$, we obtain the isomorphism $\pi_{s}^{\ast}K_{X}^{n}\simeq K_{Z_{s}}$. Hence, there exists a natural isomorphism
@@
In this way, we obtain a singular flat metric $h_{\pi_{s}^{\ast}K_{X}^{(n-1)/2}}$ as in Lemma 3.2. Because $h_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}$ is compatible with the non-degenerate symmetric bilinear product, we obtain the claim of Proposition 3.4. (See [23, Lemma 2.7].)
@@ 3.1.4 Some representation
Let $L$ be a holomorphic line bundle on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. We obtain the holomorphic line bundle $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$ of degree 00. It is equipped with the unique unitary flat connection $\nabla_{\widehat{L}}$ whose $(0,1)$-part equals the holomorphic structure of $\widehat{L}$. The monodromy representation of $\nabla_{\widehat{L}}$ induces a homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$.
@@
Let $(K_{X}^{1/2})^{\prime}$ be another square root of $K_{X}$. We obtain the holomorphic line bundle $M=(K_{X}^{1/2})^{-1}\otimes(K_{X}^{1/2})^{\prime}$ of degree 00 on $X$. It is equipped with a unitary flat connection $\nabla_{M}$ which induces $\kappa_{M}:H_{1}(X,{\mathbb{Z}})\to\{\pm 1\}$. It induces $\pi_{s}^{\ast}(\kappa_{M}):H_{1}(Z_{s},{\mathbb{Z}})\to\{\pm 1\}$. Let $\widehat{\kappa}^{\prime}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$ be the homomorphism obtained from $L$ and $(K_{X}^{1/2})^{\prime}$. Then, we have $\widehat{\kappa}_{L}=\widehat{\kappa}^{\prime}_{L}\otimes\pi_{s}^{\ast}(\kappa_{M})$.
@@ 3.2.1 Horizontal family of line bundles
Let $B$ be a simply connected Hausdorff space with a base point $b_{0}$. Let $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ be a continuous map. We set $Z_{b}:=Z_{\varphi(b)}$. Let $L_{b}$ $(b\in B)$ be holomorphic line bundles on $Z_{b}$ of degree $n(n-1)(g_{X}-1)$. For any $b_{1},b_{2}\in B$, we choose a continuous path $\gamma$ in $B$ such that $\gamma(0)=b_{1}$ and $\gamma(1)=b_{2}$, and then we obtain the isomorphisms $H_{1}(Z_{b_{1}},{\mathbb{Z}})\simeq H_{1}(Z_{b_{2}},{\mathbb{Z}})$ along the path $\varphi\circ\gamma$, which is independent of the choice of $\gamma$.
@@
Let $L$ be any holomorphic line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$. There exists a horizontal family $L_{b}$ $(b\in B)$ such that $L_{b_{0}}\simeq L$. If $L_{b}^{\prime}$ $(b\in B)$ is another horizontal family such that $L^{\prime}_{b_{0}}\simeq L$, then there exists $L_{b}\simeq L_{b}^{\prime}$ for each $b\in B$.
Proof   We set $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$, which is a holomorphic line bundle of degree 00. We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$. It induces homomorphisms $H_{1}(Z_{b},{\mathbb{Z}})\to U(1)$ and unitary flat line bundles $(\widehat{L}_{b},\nabla_{\widehat{L}_{b}})$ on $Z_{b}$ $(b\in B)$. Let $\widehat{L}_{b}$ denote the underlying holomorphic line bundle. We set $L_{b}=\widehat{L}_{b}\otimes\pi_{\varphi(b)}^{\ast}(K_{X}^{(n-1)/2})$. Then, $L_{b}$ $(b\in B)$ is a horizontal family.
@@ 3.2.2 Complex analytic property of horizontal families
Proof   We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$ as in the proof of Lemma 3.8. Because $Z_{B}$ is homotopy equivalent to $Z_{b_{0}}$, we obtain $\widehat{\kappa}_{L}^{\prime}:H_{1}(Z_{B},{\mathbb{Z}})\to U(1)$. There exists a unitary flat line bundle $(\widehat{\mathcal{L}},\nabla_{\widehat{\mathcal{L}}})$ on $Z_{B}$ which induces the homomorphism $\widehat{\kappa}_{L}^{\prime}$. Let $q_{2}:B\times X\to X$ denote the projection. We obtain the line bundle $(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2})$ on $Z_{B}$. We set
=^$\mathcal{L}=\widehat{\mathcal{L}}\otimes(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2}).$
@@ 3.2.3 Horizontal property and symmetric pairings
For $b\in B$, let $\kappa_{b}:H_{1}((Z_{b})_{0},{\mathbb{Z}})\to U(1)$ be the representation induced by the Chern connection of $\pi_{b}^{\ast}(K_{X}^{(n-1)/2})$ and the singular metric $h_{\pi_{b}^{\ast}(K_{X}^{(n-1)/2})}$. Then, $\kappa_{b}$ are independent of $b\in B$.
@@ 3.2.5 Construction of horizontal family via symmetric pairings
Proof   We use Oka-Grauert principle (see [13, §1.1.2, Theorem B’]). There exists a topological isomorphism $\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. Because $(Z_{B})_{0,Q}$ is Stein, there exists a holomorphic isomorphism $\varphi^{\prime}_{1,Q}:\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. The restriction $\varphi^{\prime}_{1,Q|(Z_{b_{0}})_{0,Q}}$ can be regarded as a holomorphic function $\alpha:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}^{\ast}$. Both $(Z_{b_{0}})_{0,Q}$ and ${\mathbb{C}}^{\ast}$ are homotopic to $S^{1}=\bigl{\{}a\in{\mathbb{C}}\,\big{|}\,|a|=1\bigr{\}}$. Because $(Z_{B})_{0,Q}$ is Stein, we can modify $\varphi^{\prime}_{1,Q}$ such that the induced map $H_{1}((Z_{b_{0}})_{0,Q},{\mathbb{Z}})\to H_{1}({\mathbb{C}}^{\ast},{\mathbb{Z}})$ is 00. Then, there exists $\beta:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}$ such that $\alpha=\exp(\beta)$. Because $(Z_{B})_{0,Q}$ is Stein, there exists $\widetilde{\beta}:(Z_{B})_{0,Q}\to{\mathbb{C}}$ whose restriction to $(Z_{b_{0}})_{0,Q}$ equals $\beta$. Then, $\varphi_{1,Q}=\exp(-\widetilde{\beta})\varphi^{\prime}_{1,Q}$ satisfies the desired condition.
@@ 3.2.7 Complement
3.2.8 Appendix: Ramification divisor
@@ 3.3.1 Horizontal family of Higgs bundles with smooth spectral curves
Let $B$ be a simply connected Hausdorff space. Let $(E_{b},\theta_{b})$ $(b\in B)$ be a family of Higgs bundles of rank $n$ of degree 00 on $X$ such that the following holds.
@@
Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ whose spectral curve is smooth. Then, there exists a horizontal family $(E_{b},\theta_{b})$ $(b\in B)$ such that $(E_{b_{0}},\theta_{b_{0}})\simeq(E,\theta)$. For another such family $(E^{\prime}_{b},\theta^{\prime}_{b})$ $(b\in B)$, there exist isomorphisms $(E_{b},\theta_{b})\simeq(E^{\prime}_{b},\theta^{\prime}_{b})$.
@@ 3.3.2 Complex analytic property of horizontal families
Let $(E,\theta)$ be a Higgs bundle of degree 00 on $X$ such that $\Sigma_{\theta}=Z_{b_{0}}$. There exists a horizontal relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $B\times X$ such that $(\mathcal{E}_{b_{0}},\theta_{\mathcal{E},b_{0}})=(E,\theta)$. If $(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$ be another such relative Higgs bundle, then there exists an isomorphism $(\mathcal{E},\theta_{\mathcal{E}})\simeq(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$.

### v3

@@ 3.1.1 Spectral curves
We recall the spectral curves by following [1]. Let $X$ be a compact Riemann surface with genus $g_{X}\geq 2$. In this section, we prefer to use the notation $K_{X}$ to denote the sheaf of holomorphic $1$-forms because we consider $K_{X}^{j/2}$. We set
@@
Let $Z_{s}$ be the subscheme of $\mathbb{P}$ obtained as the 0-set of $\xi_{s}$. It is contained in $T^{\ast}X\subset\mathbb{P}$. Let ${\boldsymbol{A}}^{\prime}_{X,n}$ denote the subset of $s\in{\boldsymbol{A}}_{X,n}$ such that $Z_{s}$ are smooth subschemes. Because $\pi_{\mathbb{P}}^{\ast}K_{X}^{n}\otimes\mathcal{O}_{\mathbb{P}}(n)$ is an ample line bundle on $\mathbb{P}$, the theorem of Bertini implies that ${\boldsymbol{A}}^{\prime}_{X,n}$ is a non-empty Zariski open subset of ${\boldsymbol{A}}_{X,n}$, and that $Z_{s}$ is irreducible for any $s\in{\boldsymbol{A}}^{\prime}_{X,n}$. Let $\pi_{s}:Z_{s}\to X$ denote the projection. For each $Q\in Z_{s}$, let $r_{Q}$ denote the ramification index of $\pi_{s}$ at $Q$. We obtain the effective divisor $\widetilde{D}_{s}=\sum(r_{Q}-1)Q$. We have $K_{Z_{s}}=\pi^{\ast}(K_{X})(\widetilde{D}_{s})$. We recall the following (see [1, §3]).
@@
For any $s\in{\boldsymbol{A}}^{\prime}_{X,n}$, the genus of $Z_{s}$ is $n^{2}(g_{X}-1)+1$. It implies $\deg(\widetilde{D}_{s})=2n(n-1)(g_{X}-1)$.
@@ 3.1.2 Singular flat metrics
Let $L_{i}$ $(i=1,2)$ be holomorphic line bundles on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. Let $U_{j}$ $(j=1,\ldots,m)$ be simply connected open subsets of $Z_{s}$ such that $U_{j}\cap U_{k}=\emptyset$ $(j\neq k)$ and that $|\widetilde{D}_{s}|\subset\bigcup U_{j}$. Let
@@ 3.1.3 A special case
The image of $\kappa_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}:H_{1}(Z_{s}\setminus|\widetilde{D}_{s}|,{\mathbb{Z}})\to U(1)$ is contained in $\{\pm 1\}$.
Proof  Recall that $Z_{s}$ is the 0-set of $\xi_{s}\in H^{0}(\mathbb{P},\pi_{\mathbb{P}}^{\ast}K_{X}^{n}\otimes\mathcal{O}_{\mathbb{P}}(n))$, and that the line bundle $\mathcal{O}_{\mathbb{P}}(n)$ is trivialized by the section $\pi_{\mathbb{P}}^{\ast}(y)$ on a neighbourhood of $Z_{s}$. (See §3.1.1 for the section $y$.) Hence, there exists a natural isomorphism between $\pi_{s}^{\ast}K_{X}^{n}$ and the normal bundle of $Z_{s}\subset T^{\ast}X$. By the symplectic structure of $T^{\ast}X$, we obtain the isomorphism $\pi_{s}^{\ast}K_{X}^{n}\simeq K_{Z_{s}}$. Hence, there exists a natural isomorphism
@@
In this way, we obtain a singular flat metric $h_{\pi_{s}^{\ast}K_{X}^{(n-1)/2}}$ as in Lemma 3.2. Because $h_{\pi_{s}^{\ast}(K_{X}^{(n-1)/2})}$ is compatible with the non-degenerate symmetric bilinear product, we obtain the claim of Proposition 3.4. (See [24, Lemma 2.7].)
@@ 3.1.4 Some representation
Let $L$ be a holomorphic line bundle on $Z_{s}$ of degree $n(n-1)(g_{X}-1)$. We obtain the holomorphic line bundle $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$ of degree 0. It is equipped with the unique unitary flat connection $\nabla_{\widehat{L}}$ whose $(0,1)$-part equals the holomorphic structure of $\widehat{L}$. The monodromy representation of $\nabla_{\widehat{L}}$ induces a homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$.
@@
Let $(K_{X}^{1/2})^{\prime}$ be another square root of $K_{X}$. We obtain the holomorphic line bundle $M=(K_{X}^{1/2})^{-1}\otimes(K_{X}^{1/2})^{\prime}$ of degree 0 on $X$. It is equipped with a unitary flat connection $\nabla_{M}$ which induces $\kappa_{M}:H_{1}(X,{\mathbb{Z}})\to\{\pm 1\}$. It induces $\pi_{s}^{\ast}(\kappa_{M}):H_{1}(Z_{s},{\mathbb{Z}})\to\{\pm 1\}$. Let $\widehat{\kappa}^{\prime}_{L}:H_{1}(Z_{s},{\mathbb{Z}})\to U(1)$ be the homomorphism obtained from $L$ and $(K_{X}^{1/2})^{\prime}$. Then, we have $\widehat{\kappa}_{L}=\widehat{\kappa}^{\prime}_{L}\otimes\pi_{s}^{\ast}(\kappa_{M})$.
@@ 3.2.1 Horizontal family of line bundles
Let $B$ be a simply connected Hausdorff space with a base point $b_{0}$. Let $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ be a continuous map. We set $Z_{b}:=Z_{\varphi(b)}$ for any $b\in B$. Let $L_{b}$ $(b\in B)$ be holomorphic line bundles on $Z_{b}$ of degree $n(n-1)(g_{X}-1)$. For any $b_{1},b_{2}\in B$, we choose a continuous path $\gamma$ in $B$ such that $\gamma(0)=b_{1}$ and $\gamma(1)=b_{2}$, and then we obtain the isomorphisms $H_{1}(Z_{b_{1}},{\mathbb{Z}})\simeq H_{1}(Z_{b_{2}},{\mathbb{Z}})$ along the path $\varphi\circ\gamma$, which is independent of the choice of $\gamma$.
@@
Let $L$ be any holomorphic line bundle on $Z_{b_{0}}$ of degree $n(n-1)(g_{X}-1)$. There exists a horizontal family $L_{b}$ $(b\in B)$ such that $L_{b_{0}}\simeq L$. If $L_{b}^{\prime}$ $(b\in B)$ is another horizontal family such that $L^{\prime}_{b_{0}}\simeq L$, then there exists an isomorphism $L_{b}\simeq L_{b}^{\prime}$ for each $b\in B$.
Proof  We set $\widehat{L}=L\otimes\pi_{s}^{\ast}(K_{X}^{(n-1)/2})^{-1}$, which is a holomorphic line bundle of degree 0. We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$. It induces homomorphisms $H_{1}(Z_{b},{\mathbb{Z}})\to U(1)$ and unitary flat line bundles $(\widehat{L}_{b},\nabla_{\widehat{L}_{b}})$ on $Z_{b}$ $(b\in B)$. Let $\widehat{L}_{b}$ denote the underlying holomorphic line bundle. We set $L_{b}=\widehat{L}_{b}\otimes\pi_{\varphi(b)}^{\ast}(K_{X}^{(n-1)/2})$. Then, $L_{b}$ $(b\in B)$ is a horizontal family.
@@ 3.2.2 Complex analytic property of horizontal families
Proof  We obtain the homomorphism $\widehat{\kappa}_{L}:H_{1}(Z_{b_{0}},{\mathbb{Z}})\to U(1)$ as in the proof of Lemma 3.8. Because $Z_{B}$ is homotopy equivalent to $Z_{b_{0}}$, we obtain $\widehat{\kappa}_{L}^{\prime}:H_{1}(Z_{B},{\mathbb{Z}})\to U(1)$. There exists a unitary flat line bundle (^,^)($\widehat{\mathcal{L}}$$\nabla_{\widehat{\mathcal{L}}}$) on $Z_{B}$ which induces the homomorphism $\widehat{\kappa}_{L}^{\prime}$. Let $q_{2}:B\times X\to X$ denote the projection. We obtain the line bundle $(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2})$ on $Z_{B}$. We set
$\mathcal{L}=\widehat{\mathcal{L}}\otimes(q_{2}\circ\pi_{B})^{\ast}(K_{X}^{(n-1)/2}).$
@@ 3.2.3 Horizontal property and symmetric pairings
For $b\in B$, let $\kappa_{b}:H_{1}((Z_{b})_{0},{\mathbb{Z}})\to U(1)$ be the representation induced by the Chern connection of $\pi_{b}^{\ast}(K_{X}^{(n-1)/2})$ with the singular metric $h_{\pi_{b}^{\ast}(K_{X}^{(n-1)/2})}$. Then, $\kappa_{b}$ are independent of $b\in B$.
@@ 3.2.5 Construction of horizontal family via symmetric pairings
Proof  We use Oka-Grauert principle (see [13, §1.1.2, Theorem B’]). There exists a topological isomorphism $\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. Because $(Z_{B})_{0,Q}$ is Stein, there exists a holomorphic isomorphism $\varphi^{\prime}_{1,Q}:\mathcal{L}_{0|(Z_{B})_{0,Q}}\simeq\mathcal{L}_{Q|(Z_{B})_{0,Q}}$. The restriction $\varphi^{\prime}_{1,Q|(Z_{b_{0}})_{0,Q}}$ can be regarded as a holomorphic function $\alpha:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}^{\ast}$. Both $(Z_{b_{0}})_{0,Q}$ and $\mathbb{C}$$^{\ast}$ are homotopic to $S^{1}=\bigl\{a\in{\mathbb{C}}\,\big|\,|a|=1\bigr\}$. Because $(Z_{B})_{0,Q}$ is Stein, we can modify $\varphi^{\prime}_{1,Q}$ such that the induced map $H_{1}((Z_{b_{0}})_{0,Q},{\mathbb{Z}})\to H_{1}({\mathbb{C}}^{\ast},{\mathbb{Z}})$ is 0. Then, there exists $\beta:(Z_{b_{0}})_{0,Q}\to{\mathbb{C}}$ such that $\alpha=\exp(\beta)$. Because $(Z_{B})_{0,Q}$ is Stein, there exists $\widetilde{\beta}:(Z_{B})_{0,Q}\to{\mathbb{C}}$ whose restriction to $(Z_{b_{0}})_{0,Q}$ equals $\beta$. Then, $\varphi_{1,Q}=\exp(-\widetilde{\beta})\varphi^{\prime}_{1,Q}$ satisfies the desired condition.
@@ 3.2.7 Complement
3.2.8 Appendix: Ramification divisors
@@ 3.3.1 Horizontal family of Higgs bundles with smooth spectral curves
Let $B$ be a simply connected Hausdorff space. Let $(E_{b},\theta_{b})$ $(b\in B)$ be a family of Higgs bundles of rank $n$ of degree 0 on $X$ such that the following holds.
@@
Let $(E,\theta)$ be a Higgs bundle of degree 0 on $X$ whose spectral curve is smooth. Let $\varphi:B\to{\boldsymbol{A}}^{\prime}_{X,n}$ is a continuous map. Then, there exists a horizontal family $(E_{b},\theta_{b})$ $(b\in B)$ such that $(E_{b_{0}},\theta_{b_{0}})\simeq(E,\theta)$ and $\Sigma_{\theta_{b}}=Z_{\varphi(b)}$. For another such family $(E^{\prime}_{b},\theta^{\prime}_{b})$ $(b\in B)$, there exist isomorphisms $(E_{b},\theta_{b})\simeq(E^{\prime}_{b},\theta^{\prime}_{b})$.
@@ 3.3.2 Complex analytic property of horizontal families
Let $(E,\theta)$ be a Higgs bundle of degree 0 on $X$ such that $\Sigma_{\theta}=Z_{b_{0}}$. There exists a horizontal relative Higgs bundle $(\mathcal{E},\theta_{\mathcal{E}})$ on $B\times X$ such that $(\mathcal{E}_{b_{0}},\theta_{\mathcal{E},b_{0}})=(E,\theta)$ and that $\Sigma_{\mathcal{E},\theta}=Z_{B}$. If $(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$ be another such relative Higgs bundle, then there exists an isomorphism $(\mathcal{E},\theta_{\mathcal{E}})\simeq(\mathcal{E}^{\prime},\theta_{\mathcal{E}^{\prime}})$.